OpenStudy (anonymous):

Sorry, that should be 2^(5/2) - 2^(3/2) =

OpenStudy (anonymous):

Even so.. \[2^\frac{5}{2} = 2^\frac{4}{2} \cdot 2^\frac{1}{2}\]